According to one survey in India, 75% of Instagram users love REELS. Suppose that 25 Instagram users (randomly selected) have been approached in the university located in vile Parle. They have been asked about their status of like/ dislike the Instagram- REELS.

a) What is the probability that exactly 15 of them would agree with the claim ( or said they love Insta-REELS) ?

b) What is the probability that Exactly 20 of them would agree with the claim (or said they love Insta-REELS) ?


1
Expert's answer
2022-03-30T12:23:06-0400

"X-number\,\,of\,\,those\,\,who\,\,like\\X\sim Bin\left( 25,0.75 \right) \\a:P\left( X=15 \right) =C_{25}^{15}\cdot 0.75^{15}\cdot 0.25^{10}=0.0416584\\b:P\left( X=20 \right) =C_{25}^{20}\cdot 0.75^{20}\cdot 0.25^5=0.164538"